# UniMoMo: Unified Generative Modeling of 3D Molecules for De Novo Binder Design

### v1

@@ Appendix C Algorithms for Training and Sampling with the Latent Diffusion Model
Appendix D Implementation Details
D.1 Baselines
@@ D.2 UniMoMo
Appendix E Implementation of Amino Acid Recovery (AAR)
@@
Appendix F Descriptions of Metrics in CBGBench
@@
Appendix G Results of Interaction Analysis for Small Molecule design
Table 9 presents the interaction analysis including Vina scores and interaction distributions for small molecule design.
Model Vina Score Vina Min Vina Dock PLIP Interaction Rank
E IMP (%) E IMP (%) E IMP (%) MPBG (%) LBE JSDOA MAEOA JSDPP MAEPP
[…]
Appendix H Case Study
In Figure 4 and 5, we provide additional designs of peptides, antibodies, and small molecules on four targets: SARS receptor binding domain, cluster of differentiation 38, Influenza H5 HA head domain, and HIV envelope glycoprotein gp160.
@@ Appendix H Case Study
Appendix I Ablations on the Number of Iterations
$N$ Peptide Antibody Small Molecule
JSDbb JSDsc JSDbb JSDsc JSDBL JSDBA
1 0.393 0.358 0.376 0.374 0.4216 0.5328
3 0.281 0.232 0.277 0.260 0.3625 0.4976
5 0.273 0.203 0.270 0.242 0.3470 0.4724
10 0.205 0.180 0.224 0.221 0.3223 0.3848

### v2

@@ Appendix D Algorithms for Training and Sampling with the Latent Diffusion Model
Appendix E Discussion on E(3)-Equivariance
We briefly outline how E(3)-equivariance is preserved throughout our framework, building upon established theoretical foundations.
For the diffusion component, GeoDiff (Xu et al., 2022) demonstrates that E(3)-equivariance is maintained via an equivariant denoising kernel (Proposition 1), which in our paper is instantiated using an equivariant transformer (Jiao et al., 2024). This transformer operates on scalar features, preserving equivariance through inner products and unbiased linear transformations applied to velocity vectors, in accordance with the scalar-based design principles Villar et al. (2021).
For the variational autoencoder, the encoder outputs latent variables using the same equivariant transformer, ensuring E(3)-equivariance by design as the latents are the direct output of the transformer. The decoder adopts a short flow matching approach, with a similar theoretical foundation as GeoDiff on equivariant flow matching (Song et al., 2023) (Theorem 4.1). Since the vector field is predicted by an E(3)-equivariant network, the decoding process consistently maintains E(3)-equivariance across all stages.
Appendix F Implementation Details
F.1 Baselines
@@ F.2 UniMoMo
Appendix G Physical Corrections on Generation
To enhance the physical plausibility of generated structures, different constraints could be incorporated into the proposed framework. Here we illustrate two common strategies to correct the clashes and the local geometries.
G.1 Clashes Avoidance
Clashes frequently occur in structures generated by deep learning models. In UniMoMo, we decouple coarse global arrangement and fine-grained local geometry using latent diffusion and a full-atom variational autoencoder, which allows us to explicitly handle steric clashes during decoding by adding repulsive forces to the predicted vector field (Eq.5) during iterative reconstruction. At each step, we compute pairwise distances between generated atoms and those in the contexts (i.e. the binding site), and apply a force when their van der Waals radii overlap beyond a threshold $\delta$ (Ramachandran et al., 2011), which we use 0.3Å in this paper. Specifically, if $r_{i}$ and $r_{j}$ are the radii of atoms $i$ and $j$, and $\vec{x}_{i}^{t}$ and $\vec{x}_{j}^{t}$ are their coordinates at time step $t$, the magnitude of the repulsive force is defined as:
$f_{ij}$ =$=\left\{\begin{array}[]{ll}r_{i}+r_{j}-\delta-\|\vec{x}_{i}^{t}-\vec{x}_{j}^{t}\|,&\|\vec{x}_{i}^{t}-\vec{x}_{j}^{t}\|<r_{i}+r_{j}-\delta,\\ 0,&else,\end{array}\right.$ (19)
Then for each iteration during decoding, the clashes can be repaired by further updating the atomic coordinates as $\vec{x}_{i}^{t}=\vec{x}_{i}^{t}+\sum_{j}\frac{\vec{x}_{i}^{t}-\vec{x}_{j}^{t}}{\|\vec{x}_{i}^{t}-\vec{x}_{j}^{t}\|}\cdot f_{ij}$, where $j$ iterates over all atoms in the fixed context.
G.2 Bond Validity
Existing literature often relies on OpenBabel (O’Boyle et al., 2011) to assign chemical bonds based on pairwise atomic distances. In contrast, our unified framework determines bonding through block-defined intra-block bonds and model-predicted inter-block bonds. While this model-driven approach allows greater flexibility, such as accommodating non-standard amino acids for future extension, it may introduce issues like invalid valencies and inconsistencies between predicted 2D topology and 3D geometry (e.g., excessively long bond lengths or distorted bond angles). To address valency violations, we rank predicted bonds by their output probability and add them sequentially. Any bond that exceeds valency constraints is discarded. For 2D–3D consistency, we compute empirical distributions of bond lengths and angles, with bond lengths discretized from 1.1Å to 1.7Å (in 0.005Å bins), and angles from $0^{\circ}$ to $180^{\circ}$ (in $2^{\circ}$ bins). If over 10% of the bonds in a generated molecule fall into zero-probability regions of these distributions, the molecule is considered 2D-3D inconsistent and discarded.
Appendix H Ablations on the Number of Iterations
$N$ Peptide Antibody Small Molecule
JSDbb JSDsc JSDbb JSDsc JSDBL JSDBA
1 0.393 0.358 0.376 0.374 0.4216 0.5328
3 0.281 0.232 0.277 0.260 0.3625 0.4976
5 0.273 0.203 0.270 0.242 0.3470 0.4724
10 0.205 0.180 0.224 0.221 0.3223 0.3848
Appendix I Implementation of Amino Acid Recovery (AAR)
@@
Appendix J Descriptions of Metrics in CBGBench
@@
Appendix K Case Study
In Figure 5 and 6, we provide additional designs of peptides, antibodies, and small molecules on four targets: SARS receptor binding domain, cluster of differentiation 38, Influenza H5 HA head domain, and HIV envelope glycoprotein gp160.
@@
Appendix L Self-Adaptive Molecular Sizes
We investigate how the choices of number of blocks influence generations on the same binding site. Using the GPCR case from § 4.4, we evaluate different block configurations averaged over 100 generations. For peptides, we vary the number of blocks into three categories: small (4–10 blocks), medium (11–17 blocks), and large (18–24 blocks). For small molecules, the original algorithm follows the standard protocol of sampling the number of blocks based on the spatial size of the binding site (Guan et al., 2023a). We modify this by shifting the distribution ±$\pm 5$ blocks to assign less or more number of blocks. As shown in Table 12, peptide results align with expectations, where larger binders generally yield lower binding energies by forming more interactions. Remarkably, for small molecules, the model exhibits a self-adaptive behavior. When fewer blocks are used, it generates larger fragments with more atoms to fill the pocket; conversely, with more blocks, it opts for smaller fragments to avoid steric clashes, indicating context-aware flexibility in generation.
Table 12: Comparisons of energies and number of atoms in each block for peptides and small molecules under different settings of designated number of blocks.
Peptide Small Molecule
#Blocks Rosetta dG #Blocks Vina Score (dock) Avg. #Atoms per Block
small (4-10) -8.09 small (n-5) -6.72 5.10
medium (11-17) -16.93 medium (n) -7.42 4.06
big (18-24) -22.39 big (n+5) -7.44 3.35